OpenStudy (anonymous):

If cosx=−23 and sinx<0, findsinx

OpenStudy (anonymous):

A) −95 B)−95 C)−35 D)−13 E)−914

OpenStudy (anonymous):

this is what you do if you want easy way. look up "half angle formula" for sine, which says √\[ \sin(x2)=+-\sqrt{?}1 - \cos(x)2 /2\] replace replace cos(x) by -23